# Improving criminal case management through Machine Learning system

### v1

@@ Abstract
Keywords machine learning $\cdot$ criminal justice $\cdot$ case prioritization $\cdot$ prosecutorial backlog $\cdot$ public sector AI $\cdot$ decision support systems $\cdot$ Mexico $\cdot$ Random Forest
@@ 1 Introduction
The prosecutorial backlog in Mexico’s criminal justice system poses a major institutional challenge. Across the country, prosecutors’ offices are responsible for overseeing every stage of criminal proceedings—opening investigations, collecting evidence, promoting alternative solutions, and pursuing criminal action in court. However, growing caseloads and limited resources have led to large volumes of unresolved cases. Between 2018 and 2022, the number of open files nationwide increased by nearly 60%, with over 2.9 million active investigations reported by the end of 2022 MexicoEvalua2023; INEGI2023.
@@ 4 Features
Using data from the Plataforma de Integración de Expedientes (PIE), we construct a set of features designed to capture how the Prosecutor’s Office interacts with each case over time. Our unit of prediction is the individual case record, specifically those that remain open in the Módulo de Atención Temprana (MAT) at the time of analysis. The features span both static characteristics of the case and dynamic indicators of prosecutorial activity, reflecting not only the nature of the crime but also the administrative and legal progression of the case within the office. This strategy aligns with recent work emphasizing event-driven features and longitudinal context in high-stakes domains Chouldechova2018; Kleinberg2018.
@@
This feature set reflects principles from interpretable machine learning design, prioritizing temporally-grounded and institutionally meaningful variables Lakkaraju2016; Molnar2020. By focusing on longitudinal event structures rather than static case attributes, the model captures not just what kind of case is being handled, but how it is evolving—by whom, for how long, and with what apparent momentum.
@@ 5 Methodology
To simulate real-time deployment, we use a rolling temporal cross-validation strategy Roberts2017. At each prediction point, we train the model only on data available up to that moment, ensuring no information leakage into the future. This approach has been recommended for high-stakes public sector applications where real-time decision support must mirror deployment conditions berg2021validating.
Model training uses case histories from 2014 to 2024, with predictions and evaluation focused on the 2023–2024 period. We evaluate models using Precision@300 and Recall@300, which are operationally relevant metrics aligned with institutional constraints, following similar practices in ML-for-policy research lakkaraju2015interpretable; jung2017simple.
@@ 6 Results
Figure 1 presents model performance over time, showing that ensemble methods—particularly Random Forests—consistently outperform simpler baselines. This aligns with findings from other ML applications in legal and public service domains, where ensemble classifiers have proven effective at detecting latent patterns in administrative data berk2018fairness; kaminski2020predictability.
@@ 6.1 Feature Importance
6.2 Potential Prescription in Low-Priority Cases
@@ 7 Implementation and Future Steps
While the selected model shows strong performance on historical data, we need to validate its effectiveness in a real-time operational setting before integrating it into prosecutorial workflows. To understand the real-world impact of the model and the prioritization system it supports, we plan to conduct a randomized controlled trial (RCT) in the Módulo de Atención Temprana (MAT). This evaluation aims to determine whether prioritizing cases based on the model’s output leads to measurable improvements in prosecutorial throughput under routine operating conditions kleinberg2018human; jung2020eliciting.
The selected model—a Random Forest classifier—achieved an average Precision@300 of 0.74 across historical validation periods. While this indicates strong discriminatory performance, retrospective accuracy alone cannot establish real-time utility. We structure the RCT to causally evaluate whether model-informed prioritization increases the number of cases finalized or reduces time to resolution relative to the status quo berg2021validating; chouldechova2018case.
@@
We will conduct the experiment over six consecutive weeks, yielding approximately 1800 treatment cases and 1800 control cases. We will follow all cases for six months and compare outcomes along operationally relevant dimensions such as resolution rates and time to completion chouldechova2018case. This design tests the effect of prioritized visibility itself, rather than any procedural or staffing change.
@@ 8 Ethical Considerations and Limitations
However, any model trained on historical administrative data may reflect patterns of past bias or structural inequality. For example, the prioritization logic may implicitly favor case types, jurisdictions, or procedural pathways that have historically moved through the system more quickly. This could unintentionally reinforce disparities if not carefully monitored Barabas2018; kleinberg2018human.

### v2

# Designing AI for Prosecutorial Governance: Case Prioritization and Statutory Oversight in Mexico

@@ Abstract
Keywords artificial intelligence $\cdot$ digital government $\cdot$ criminal justice $\cdot$ algorithmic governance $\cdot$ case prioritization $\cdot$ public sector AI $\cdot$ decision support systems $\cdot$ Mexico
@@ 1 Introduction
Digital-era governance has fundamentally transformed how public institutions manage information, deliver services, and exercise administrative discretion (Dunleavy et al., 2006; Fountain, 2001). In criminal justice systems, where decisions carry profound consequences for individuals and communities, this transformation presents both opportunities and challenges. Algorithmic systems can help institutions process large caseloads more efficiently, but they also raise concerns about transparency, accountability, and the preservation of professional judgment (Janssen and Kuk, 2016; Veale et al., 2018). As governments increasingly deploy artificial intelligence (AI) in high-stakes domains, understanding how to design, implement, and oversee these systems becomes essential for both researchers and practitioners (Wirtz et al., 2019; Bullock et al., 2020).
This paper contributes to the growing literature on AI in government by presenting a case study of algorithmic decision support in a Mexican state prosecutor’s office. We describe the design, implementation, and governance of a machine learning system developed to address two interrelated institutional challenges: (1) supporting prosecutors in identifying cases likely to be resolved in the near term, and (2) providing systematic oversight of cases that may have exceeded statutory time limits. The system thus serves a dual function—enhancing operational efficiency while simultaneously enabling legal accountability. This dual-purpose design reflects a broader lesson for public sector AI: algorithmic tools need not be limited to efficiency gains but can also be structured to support institutional oversight and democratic governance.
Our analysis responds to calls for empirical research on AI implementation in government contexts outside the Global North. While much of the existing literature on algorithmic governance draws on cases from the United States and Western Europe, criminal justice institutions in Latin America face distinct challenges related to institutional capacity, digital infrastructure, and legal frameworks. By documenting the design choices and implementation experience in a Mexican prosecutor’s office, we offer insights that may inform similar efforts in other resource-constrained settings.
The prosecutorial backlog in Mexico’s criminal justice system poses a major institutional challenge. Across the country, prosecutors’ offices are responsible for overseeing every stage of criminal proceedings—opening investigations, collecting evidence, promoting alternative solutions, and pursuing criminal action in court. However, growing caseloads and limited resources have led to large volumes of unresolved cases. Between 2018 and 2022, the number of open files nationwide increased by nearly 60%, with over 2.9 million active investigations reported by the end of 2022 México Evalúa (2023); Instituto Nacional de Estadística y Geografía (INEGI) (2023).
@@ 4 Features
Using data from the Plataforma de Integración de Expedientes (PIE), we construct a set of features designed to capture how the Prosecutor’s Office interacts with each case over time. Our unit of prediction is the individual case record, specifically those that remain open in the Módulo de Atención Temprana (MAT) at the time of analysis. The features span both static characteristics of the case and dynamic indicators of prosecutorial activity, reflecting not only the nature of the crime but also the administrative and legal progression of the case within the office. This strategy aligns with recent work emphasizing event-driven features and longitudinal context in high-stakes domains Chouldechova et al. (2018); Kleinberg et al. (2018).
@@
This feature set reflects principles from interpretable machine learning design, prioritizing temporally-grounded and institutionally meaningful variables Lakkaraju et al. (2016); Molnar (2020). By focusing on longitudinal event structures rather than static case attributes, the model captures not just what kind of case is being handled, but how it is evolving—by whom, for how long, and with what apparent momentum.
@@ 5 Methodology
To simulate real-time deployment, we use a rolling temporal cross-validation strategy Roberts et al. (2017). At each prediction point, we train the model only on data available up to that moment, ensuring no information leakage into the future. This approach has been recommended for high-stakes public sector applications where real-time decision support must mirror deployment conditions.
Model training uses case histories from 2014 to 2024, with predictions and evaluation focused on the 2023–2024 period. We evaluate models using Precision@300 and Recall@300, which are operationally relevant metrics aligned with institutional constraints, following similar practices in ML-for-policy research Lakkaraju et al. (2016); Jung et al. (2017).
@@
A key design principle underlying this system is that algorithmic tools in government should serve multiple governance functions when possible. Rather than building separate systems for productivity enhancement and legal oversight, we intentionally designed a unified framework that addresses both needs from a single prediction model. The same probability estimates that identify cases likely to be resolved in the near term also reveal, at the opposite tail of the distribution, cases that appear stalled and may warrant administrative review. This dual-purpose architecture reflects an emerging principle in public sector AI: that well-designed algorithmic systems can simultaneously support operational efficiency and institutional accountability (Peeters and Widlak, 2018; Veale et al., 2018).
@@ 6 Results
Figure 1 presents model performance over time, showing that ensemble methods—particularly Random Forests—consistently outperform simpler baselines. This aligns with findings from other ML applications in legal and public service domains, where ensemble classifiers have proven effective at detecting latent patterns in administrative data Berk et al. (2018); Wang et al. (2020).
@@ 6.1 Feature Importance
6.2 Oversight Function: Identifying Potentially Prescribed Cases
@@ 7 Implementation and Future Steps
While the selected model shows strong performance on historical data, we need to validate its effectiveness in a real-time operational setting before integrating it into prosecutorial workflows. To understand the real-world impact of the model and the prioritization system it supports, we plan to conduct a randomized controlled trial (RCT) in the Módulo de Atención Temprana (MAT). This evaluation aims to determine whether prioritizing cases based on the model’s output leads to measurable improvements in prosecutorial throughput under routine operating conditions Kleinberg et al. (2018).
The selected model—a Random Forest classifier—achieved an average Precision@300 of 0.74 across historical validation periods. While this indicates strong discriminatory performance, retrospective accuracy alone cannot establish real-time utility. We structure the RCT to causally evaluate whether model-informed prioritization increases the number of cases finalized or reduces time to resolution relative to the status quo Chouldechova et al. (2018).
@@
We will conduct the experiment over six consecutive weeks, yielding approximately 1800 treatment cases and 1800 control cases. We will follow all cases for six months and compare outcomes along operationally relevant dimensions such as resolution rates and time to completion Chouldechova et al. (2018). This design tests the effect of prioritized visibility itself, rather than any procedural or staffing change.
@@
8 Discussion
This case study offers several insights for the design and governance of AI systems in public sector contexts. First, the dual-purpose architecture demonstrates that algorithmic tools can serve both operational efficiency and institutional oversight objectives simultaneously. By designing a single prediction framework that supports near-term prioritization at one tail of the distribution and statutory review at the other, we avoided the need for separate systems while ensuring that both governance functions receive attention. This approach aligns with calls for AI systems that embed accountability mechanisms from the design stage rather than treating oversight as an afterthought (Veale et al., 2018; Peeters and Widlak, 2018).
Second, the implementation experience highlights the importance of institutional fit. The system was designed around existing workflows rather than requiring organizational restructuring. Weekly lists are delivered through channels already familiar to prosecutors, and the ranked format aligns with how staff already conceptualize case prioritization. This “low-overhead” approach, which aims to enhance rather than replace professional judgment, appears essential for adoption in resource-constrained settings where staff have limited capacity to learn new systems or adapt to changed procedures.
Third, the case illustrates how administrative data infrastructures shape the possibilities for algorithmic governance. The Plataforma de Integración de Expedientes (PIE) provides the digital foundation that makes this kind of analysis possible, but its structure also constrains what can be predicted. The system captures procedural events but not the substantive content of case files; it records timestamps but with some noise; it classifies crimes at a level of aggregation that introduces uncertainty into prescription calculations. These limitations are not unique to Zacatecas—they reflect the broader reality that algorithmic systems in government must work with administrative data designed for other purposes (Janssen and Kuk, 2016).
Finally, this work contributes to the emerging literature on AI in government contexts outside the Global North. The challenges faced by the Zacatecas prosecutor’s office—high caseloads, limited staff, persistent backlogs—are common across Latin American criminal justice systems and in many other developing country contexts. By documenting both the potential and the limitations of algorithmic decision support in this setting, we hope to inform future efforts to design AI systems that are appropriate for the institutional realities of diverse governance contexts.
@@ 9 Ethical Considerations and Limitations
However, any model trained on historical administrative data may reflect patterns of past bias or structural inequality. For example, the prioritization logic may implicitly favor case types, jurisdictions, or procedural pathways that have historically moved through the system more quickly. This could unintentionally reinforce disparities if not carefully monitored Barabas et al. (2018); Kleinberg et al. (2018).
@@
Data Availability
The data used in this study contains sensitive criminal justice information and cannot be publicly shared due to legal and privacy restrictions under Mexican law. The administrative records from the Plataforma de Integración de Expedientes (PIE) include personal identifiers, crime classifications, and case disposition information that are protected under federal and state data protection regulations. Aggregated results, model specifications, and additional methodological details are available from the corresponding author upon reasonable request.
Declaration of Competing Interests
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Funding
This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors. The collaboration between Tecnológico de Monterrey and the Fiscalía General de Justicia del Estado de Zacatecas was conducted as part of ongoing institutional cooperation on public sector innovation.
Author Contributions (CRediT)
Fernanda Sobrino: Methodology, Software, Formal analysis, Writing – original draft, Visualization. Adolfo De Unánue: Conceptualization, Methodology, Supervision, Writing – review & editing. Edgar Hernández: Data curation, Software. Patricia Villa: Investigation. Elena Villalobos: Visualization, Data curation. David Aké: Project administration. Stephany Cisneros: Investigation. Cristian Paul Camacho Osnay: Conceptualization, Resources, Project administration. Armando García Neri: Conceptualization, Supervision, Resources. Israel Hernández: Data curation, Resources.